OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1 + sec^2x sin^2x = sec^2x

OpenStudy (anonymous):

do you know your Pythagorean identities? you know that sec^2x will make a 1+tan^2x and evaluate the left side by writing the trig identities in variables. is how I did it. if you need me to explain more I would like an attempt from you

OpenStudy (anonymous):

Yeah, an explanation would help. I'm confused.

OpenStudy (anonymous):

I will show you what I have so far

OpenStudy (anonymous):

alright. not my comment that sec^2x =1+tan^2x now write sec^2x * sin^2x as variables where \[\frac{ h^2 }{ a^2 }\] is secant ^2x then write out sin^2x the same way and look at them. should be very obvious

OpenStudy (anonymous):

*note my comment

OpenStudy (anonymous):

\[1 + \sec^2x \sin^2x\] \[1 + (1+ \tan^2x)(1-\cos^2x)\] This is how far I got. I don't know what you mean by h^2/a^2

OpenStudy (anonymous):

you missing my point \[1+\frac{ h^2 }{ a^2 }*\frac{ o^2 }{ h^2 }=\sec^2x\] where \[\sec^2x=1+\tan^2x\] for the right side of the equation. leave the left side as the fractions. o=opposite a=adjacent h=hypotenuse. now what is special about the left side now

OpenStudy (anonymous):

\[\frac{ h^2 }{ a^2 } * \frac{ o^2 }{ h^2 } = 1+\tan^2x\]

OpenStudy (anonymous):

now cross out your h^2 and don't forget about the 1+ on the left side and what is o/a?

OpenStudy (anonymous):

\[\tan^2x ?\]

OpenStudy (anonymous):

yes sir, why the question? this seems so fundamental

OpenStudy (anonymous):

the question mark after the tanx comment

OpenStudy (anonymous):

well, Is this writing a proof?

OpenStudy (anonymous):

not rigorously, but you just said verify the equation. This process should do jjust fine for these purposes. if you want a rigorous proof. im the wrong dude to be asking.

OpenStudy (anonymous):

Okay, that's fine. Thanks for responding! :)

OpenStudy (anonymous):

no problem.